Exercise 3.1.4: Write ay” + by + cy = f(x) as a first order system of ODES.

Please show all steps, thanks! (I know this problem has already been solved by others, but they all have different answers...)
Answer #1

Solution :- 29+ by+cy = f(n) —® put nicy me = y Then, m=y=h=y = 1=2 Me = y => M2!= from 6 ay +byt cy = f(x) say =